OpenStudy (yuii):

The ordered pairs model an exponential function, where j is the function name and e is the input variable. {(1, 10),  (2, 50) , ( 3, 250) , (4, 1250)} What is the function equation in sequence notation?

OpenStudy (welshfella):

Hint: 10*1 = 10 10*5 = 50 10*25 = 250 How can we write the second part of the multiplication as an exponential?

OpenStudy (welshfella):

1 = 5^0 5 = 5^1 25 = 5^2 and so on..

OpenStudy (welshfella):

any ideas? J = what?

OpenStudy (welshfella):

the exponent will be e as stated in the question

OpenStudy (yuii):

Is it 3125? I don't know.

OpenStudy (emogirlatemoooocow):

And here @Yuii